# MMAO-Bench: MultiModal All in One Benchmark Reveals Compositional Law between Uni-modal and Omni-modal in OmniModels

### v1

@@ 4.3.1 Ablation Visual Understanding
To evaluate the enhancement of visual information on model understanding, we sequentially assessed the performance of Audio (only auditory information), +Caption (auditory + Caption information), and +Visual (auditory information + visual information), where captions were derived from descriptions of videos by Gemini-2.5-Pro.
In the case of audio-only input, most models’ metrics are significantly lower than when both auditory and visual information are provided, and the metrics exhibit a random distribution around 25%. Gemini-2.5-Pro identified problem shortcuts through methods such as elimination, achieving a score of 47.9, which will be optimized in subsequent versions.
After providing the caption information, all model metrics showed significant increases, but the magnitude of the rise varied considerably. In terms of perceptual ability, the Gemini series generally improved by 20 points, while the Qwen-2.5-omni series and Baichuan-omni-1.5 saw an improvement of over 8 points. The MiniCPM-o-2.6 showed a smaller increase of around 4 points, whereas the Ming-lite-omni-1.5 barely improved at all. This indicates that different models have varying abilities to reconstruct image content based on textual descriptions, and similar phenomena are observed in reasoning capabilities as well.
Compared with caption information, the original visual input brings more information. On models with better baseline capabilities, such as Ming-lite-omni-1.5 and Qwen-2.5-omni-7B, caption information is far less than the original visual information. On models with slightly weaker baseline capabilities, the two are approximately the same. Interestingly, the Gemini series has a very high utilization rate of caption information, and even Gemini-2.0-Flash performs better with +caption than with +visual.
@@ 4.3.2 Ablation Audio Understanding
We categorize sound types into ambient sounds and dialogue, and provide their transcription versions based on Gemini-2.5-pro, where dialogue is given dialogue text, and ambient sounds are given sound descriptions.
Similar to visual information, in the case of only providing auditory information, most models’ metrics are significantly lower than when both auditory and visual information are provided; after providing speech transcripts, the improvement is again more pronounced in the Gemini series.
MiniCPM-o-2.6 performs better than +Audio on +ASR, confirming that the bottleneck of this model lies in audio understanding.
Table 5: Ablation of audio understanding ability.
Model Perception Reasoning Overall
Video +ASR +Audio Video +ASR +Audio Video +ASR +Audio
MiniCPM-o-2.6 27.75 31.79 27.75 27.07 29.98 26.40 27.26 30.48 26.77
Qwen-2.5-omni-3B 27.17 30.06 32.37 18.79 23.94 25.28 21.13 25.65 27.26
Baichuan-omni-1.5 26.01 27.17 31.79 27.07 23.04 29.75 26.77 24.19 30.32
Ming-lite-omni-1.5 32.37 35.84 37.57 23.04 27.52 29.98 25.65 29.84 32.10
Qwen-2.5-omni-7B 30.06 41.04 43.35 24.16 27.52 29.31 25.81 31.29 33.23
Gemini-2.0-Flash 31.79 55.49 49.71 29.08 47.43 44.97 29.84 49.68 46.29
Gemini-2.5-Flash 38.73 72.83 69.94 37.58 68.68 67.56 37.90 69.84 68.23
Gemini-2.5-Pro 46.82 76.88 81.50 47.20 75.62 74.72 47.10 75.97 76.61
In terms of environmental sound, most open-source models perform poorly. The two more capable models, Ming-lite-omni-1.5 and Qwen-2.5-omni-7B, show better performance in dialogue than in environmental sound. Additionally, we observed that the audio information of MiniCPM-o-2.6 and Ming-lite-omni-1.5 has side effects in environmental sound scenarios. When audio descriptions replace the original audio, both models show significant improvement, indicating that Ming-lite-omni-1.5 and Qwen-2.5-omni-7B have relatively weaker understanding capabilities of audio elements.
In conversational scenarios, the Gemini series achieves ASR results that are nearly on par with, or even surpass, the original audio quality. This distinctive feature is absent in most other open-source models, with the exception of MiniCPM-o-2.6, where the +ASR results exceed +Audio 4 points.
@@
Visual +Caption +Audio Visual +ASR +Audio Visual +ASR/Caption +Audio
MiniCPM-o-2.6 33.33 31.11 28.89 26.70 30.54 26.53 27.26 30.48 26.77
Qwen-2.5-omni-3B 15.56 31.11 28.89 21.64 25.31 27.23 21.13 25.65 27.26
Baichuan-omni-1.5 26.67 22.22 33.33 26.70 24.43 30.02 26.77 24.19 30.32
Ming-lite-omni-1.5 35.56 37.78 28.89 24.96 29.32 32.29 25.65 29.84 32.10
Qwen-2.5-omni-7B 28.89 40.00 26.67 25.65 30.72 33.86 25.81 31.29 33.23
Gemini-2.0-Flash 37.78 42.22 40.00 29.32 50.44 46.77 29.84 49.68 46.29
Gemini-2.5-Flash 53.33 62.22 51.11 36.47 70.51 69.63 37.90 69.84 68.23
Gemini-2.5-Pro 57.78 71.11 75.56 46.07 76.44 76.79 47.10 75.97 76.61
@@ 4.4 Benchmark Analysis
In this section, we verify the effectiveness of MMAO-Bench on three aspects, the performance of multi-step open-ended question, the performance of dataset compression and the benchmark comparison with other open-sourced benchmarks.
@@ 4.4.1 Multi-Step Open-Ended Question Analysis
In this work, we introduce a new type of evaluation method, multi-step open-ended question, which effectively assess the complex reasoning ability, especially appears in cross-modality understanding. As shown in Table.7, the experimental results demonstrate the stratified performance among different models. Gemini-2.5-Pro leads with a comprehensive score of 68.84, showcasing its outstanding capabilities as a top-tier closed-source model. Following closely are Gemini-2.5-Flash and Gemini-2.0-Flash, which also form a clear leading tier. Among open-source models, Qwen-2.5-omni-7B performs best, significantly outperforming other smaller-scale open-source models. This indicates that model scale and advanced architecture remain key factors in addressing complex multi-question tasks.
As the depth of questions increases, model performance generally declines. This is a very significant trend, also validating the effectiveness of our dataset design. For almost all models, their performance on small question 1 is better than on small question 2, while the score for small question 3+ is the lowest. For example, Gemini-2.5-Pro’s score on the "perception" task dropped from 26.60 on small question 1 to 21.79 on small question 3+; on the "reasoning" task, the score sharply decreased from 38.2 on small question 1 to 21.33 on small question 3+. This performance decay phenomenon reveals the common challenges current large models face in handling long-range dependencies, maintaining conversational context, and conducting multi-step chain-of-thought reasoning. Initial perception and simple reasoning are relatively easy, but as conversations deepen, the accumulation of errors and loss of context lead to significantly lower success rates in answering more complex subsequent questions.
Reasoning ability is the key bottleneck that distinguishes model performance. Comparing the total scores of "perception" and "reasoning" dimensions, we found that all models scored higher on perception tasks than on reasoning tasks. For example, Gemini-2.5-Pro’s perception total score is 78.4, while its reasoning total score is 66.45. Among open-source models, Qwen-2.5-omni-7B’s perception total score (40.8) is significantly higher than its reasoning total score (26.20). This indicates that although models have already acquired basic perception capabilities (such as object recognition and sound recognition), there is still significant room for improvement in their ability to transform this perceptual information into complex logical, spatial, or causal reasoning. Our dataset successfully amplifies this capability gap through its progressive question design.
@@
Baichuan-omni-3B 8.95 7.75 5.38 24.2 11.8 9.6 9.33 19.55 9.52 8.12 5.87 20.48
MiniCPM-6-2.6B 10.65 9.35 8.11 34.2 22.4 11.8 6.67 18.85 11.16 8.12 6.33 21.92
Qwen-2.5-omni-3B 10.95 8.59 8.76 34.2 19.4 13.2 5.33 23.60 12.64 9.52 8.33 25.72
Ming-lite-omni 8.35 7.20 6.29 33.2 20.8 12.4 4.67 24.30 12.68 9.96 7.69 26.08
Qwen-2.5-omni-7B 12.85 9.75 7.90 40.8 18.4 17.6 16.00 26.20 13.96 11.33 8.92 29.12
Gemini-2.0-Flash 17.55 13.60 11.32 56.2 27.8 22.0 21.33 37.15 19.6 15.28 12.56 40.96
Gemini-2.5-Flash 23.40 25.05 20.66 72.4 35.2 31.4 26.00 59.40 25.76 26.32 21.32 62.00
Gemini-2.5-Pro 26.60 29.10 21.79 78.4 38.2 33.8 21.33 66.45 28.92 30.04 21.74 68.84
@@ 4.4.2 Dataset Compression
Baseline dataset consists of  8000 samples including 15 open-source benchmarks (e.g. MathVista and MMAU, details see appendix) and 12 models evaluation results on them, which splits into 8/4 on models as training/test set. Kmeans is used to cluster with K=48. To eliminate the random factor, we conduct 5-fold settings and evaluate 10 times on each setting.
To verify the compression performance, we use metrics Margin of Error(MoE), Root Mean Square Error(RMSE), CI Coverage, Spearman’s linear correlation coefficient(SRCC) and Pearson’s linear correlation coefficient(PLCC). The experimental result is shown in Figure.8, when compression rate is 10%, the coverage is nearly 100% and PLCC is above 0.98, SRCC is above 0.98, RMSE is 0.06, which means the model ranking is consistent.
@@ 4.4.3 Benchmark Comparison
Besides dataset quality and evaluation efficiency, a good benchmark should provide sufficient metric discriminability and a progressive difficulty scale to drive model capability growth. Compared with existing omni-modal benchmarks (including OmniBench(Li et al., 2024b), WorldSense(Hong et al., 2025) and AVOdyssey(Gong et al., 2024)), MMAO-Bench has relatively consistent rankings but a higher standard deviation of 19.5 in evaluation scores, while the best other OmniBench only has 12.75, showing the MMAO-Bench has greater discriminative ability. A simple way to improve discrimination is increasing problem difficulty, which can’t guide model’s growth. As a verification, we can see that Gemini-2.5-pro scored 76.61 on MMAO-Bench, but only 65.49 on OmniBench, indicating that MMAO-Bench has appropriate difficulty levels.
@@ 5 Conclusion
1. With unified data framework in MMAO-Bench, we found that the multimodal capability may not simply be a linear superposition of uni-modalities, but rather follows a significant multiplicative relationship. The evaluation results show that it manifests as a bottleneck effect on weak models, while exhibiting synergistic promotion on strong models.
@@
2. Both uni-modal and omni-modal understanding capability of the Gemini series far surpasses existing open-source omni-models. The Gemini-2.5-Pro shows comparable perception capability with human experts but still has a performance gap in reasoning aspect.
@@
3. Besides better dataset quality and evaluation efficiency, MMAO-Bench can provide sufficient metric discriminability and a progressive difficulty scale to drive model capability growth.
@@
In the future work, we will extend the dataset’s scale by human-in-the-loop automatic pipeline and hold a private test set to avoid hacking. The ability coverage also needs to extend to more difficult reasoning tasks like STEM and code. At the same time, the relationship among cross-modals understandings and how to improve them are still exciting problem to explore.

### v2

@@ 4.3.1 Ablation Visual Understanding
To quantify the contribution of visual information, we conducted an ablation study with three settings: audio-only (Audio), audio plus high-quality textual captions of the visual scene (+ Caption), and the full audio-visual input (+ Visual). The captions were generated by Gemini-2.5-Pro to ensure descriptive richness. Results are detailed in Table.4.
With only audio input, most models’ performance drops to a level near random guessing (around 20-28%), confirming the critical role of visual context. A notable exception is Gemini-2.5-Pro, which scores 40.34, suggesting an ability to leverage linguistic cues or shortcuts within the questions even without visual data.
The introduction of Caption information yields significant but highly variable performance gains. Powerful models like the Gemini series and Qwen-3-Omni-30B-A3B demonstrate a substantial leap in performance (gains of 20-25 points), showcasing their strong ability to reconstruct scenes from textual descriptions. In contrast, models like MiniCPM-O-2.6 and Ming-lite-Omni-1.5 show minimal improvement, indicating a weaker capacity for this text-to-vision reasoning.
Comparing Caption against full Visual input reveals a fascinating dichotomy. For the most capable model, Gemini-2.5-Pro, direct visual information provides a clear advantage over captions (70.90 vs. 65.10), proving that raw visual data contains nuances that text cannot fully capture. However, for several other models, including Gemini-2.0-Flash and the powerful Qwen-3-Omni-30B-A3B, performance with captions is surprisingly on par with, or even slightly exceeds, that with direct visual input. This suggests that for these models, the language processing pathway may be more adept at extracting semantic meaning than their own visual encoders, highlighting a potential imbalance in their multimodal processing capabilities.
@@ 4.3.2 Ablation Audio Understanding
To isolate the impact of auditory information, we evaluated models under three conditions: visual-only (Visual), visual plus transcribed audio (+Caption), and the full audio-visual input (+Audio). We further divided the audio into two categories: Environment sounds, for which we provided textual descriptions, and Speech, for which we provided ASR transcripts. All transcriptions were generated by a powerful multimodal model. The results are presented in Table.5.
The Visual-only setting results in significantly lower performance across all models, with Overall scores ranging from 21.20 to 33.70. This confirms the critical role of auditory context in multimodal understanding. The introduction of textual audio descriptions (+Caption) substantially boosts performance across the board. The improvement is particularly dramatic for high-capacity models like Gemini-2.5-Pro (+31.0 points Overall) and Qwen-3-Omni-30B-A3B (+17.4 points Overall), demonstrating their strong ability to integrate textual information.
The comparison between +Caption and +Audio reveals crucial insights into the models’ raw audio processing capabilities. In environmental sound scenarios, understanding raw audio remains a significant challenge for most open-source models. For instance, Qwen-2.5-Omni-3B, MiniCPM-O-2.6, and Ming-lite-Omni-1.5 all exhibit considerably higher performance with textual descriptions (+Caption) than with the original audio (+Audio). This suggests that their audio encoders struggle to extract meaningful features from complex non-speech sounds, making them prefer clean textual summaries. In contrast, the most capable models—Gemini-2.5-Pro, Gemini-2.5-Flash, and Qwen-3-Omni-30B-A3B—demonstrate superior audio understanding by scoring higher in the +Audio setting, indicating they can extract richer information directly from the audio signal than is present in the provided caption.
In conversational (Speech) scenarios, the results are more nuanced. The top-performing Gemini-2.5-Pro shows a substantial advantage with raw audio over ASR transcripts (+Audio 72.16 vs. +Caption 66.00), indicating it effectively leverages paralinguistic cues such as tone, emotion, and prosody that are lost in transcription. Conversely, several other models, including the Qwen series and MiniCPM-O-2.6, perform slightly better with ASR transcripts (+Caption) than with raw audio. This points to a common bottleneck where imperfections in their audio encoders are a greater liability than the information lost during ASR, making clean text a more reliable input. Notably, Gemini-2.5-Flash achieves nearly identical scores in both settings, suggesting its ASR and audio understanding capabilities are exceptionally well-aligned.
@@
Visual +Caption +Audio Visual +Caption +Audio Visual +Caption +Audio
Qwen-2.5-Omni-3B 26.28 41.03 34.62 24.76 26.66 26.54 25.00 28.90 27.80
MiniCPM-O-2.6 26.92 39.74 34.62 28.08 28.44 27.49 27.90 30.20 28.60
Ming-lite-Omni-1.5 31.41 43.59 35.26 22.27 25.59 27.73 23.70 28.40 28.90
Baichuan-Omni-1.5 25.64 32.05 28.85 23.70 23.58 29.86 24.00 24.90 29.70
Qwen-2.5-Omni-7B 30.77 41.03 37.18 24.41 33.06 31.75 25.40 34.30 32.60
Qwen-3-Omni-30B-A3B 32.05 48.08 48.72 23.58 41.23 40.88 24.90 42.30 42.10
Gemini-2.0-Flash 25.00 48.08 45.51 22.87 48.93 44.79 23.20 48.80 44.90
Gemini-2.5-Flash 17.95 48.72 49.36 21.80 55.09 55.21 21.20 54.10 54.30
Gemini-2.5-Pro 32.69 57.69 64.10 33.89 66.00 72.16 33.70 64.70 70.90
@@ 4.4 Benchmark Analysis
In this section, we verify the effectiveness of UNO-Bench on three aspects, the performance of multi-step open-ended question, the performance of dataset compression and the benchmark comparison with other open-source benchmarks.
@@ 4.4.1 Multi-Step Open-Ended Question Analysis
In this work, we introduce a new type of evaluation method, multi-step open-ended question, which effectively assess the complex reasoning ability, especially appears in cross-modality understanding.
As shown in Table.6, the experimental results on our multi-step open-ended questions reveal a clear performance stratification among models. Gemini-2.5-Pro establishes itself as the top-tier model with an overall score of 57.32, with Gemini-2.5-Flash (47.08) and Gemini-2.0-Flash (38.56) forming a distinct second tier. Among open-source models, Qwen-3-Omni-30B-A3B emerges as the clear leader with a score of 37.08, significantly outperforming smaller-scale models like Qwen-2.5-Omni-7B (27.72). This starkly illustrates that both advanced architecture and model scale are pivotal factors for success in complex, multi-turn multimodal tasks.
As the depth of questions increases from Q1 to Q3+, most models exhibit a general decline in performance, confirming the effectiveness of our dataset’s progressive difficulty. For instance, the leading open-source model, Qwen-3-Omni-30B-A3B, sees its overall score drop from 18.08 on the first question (Q1) to 14.18 (Q2) and further to 11.42 (Q3+). This decay highlights a common challenge for current models in handling long-range dependencies, maintaining conversational context, and performing multi-step reasoning. However, a notable exception is Gemini-2.5-Pro, whose performance on the second question (Q2) surpasses its score on the first (24.48 vs. 23.44), before declining on subsequent questions. This unique pattern suggests a superior ability to utilize the context from the initial turn to enhance its understanding and response in the subsequent turn, a capability not observed in other models.
Reasoning ability remains the key bottleneck that differentiates model performance. For all open-source models and the lower-tier Gemini models, scores on Perception tasks are considerably higher than on Reasoning tasks. The gap is particularly pronounced for Qwen-3-Omni-30B-A3B, which scores 53.8 in Perception but only 32.9 in Reasoning. This indicates that while these models have developed solid foundational perception capabilities, converting this perceptual input into complex logical or causal reasoning remains a major hurdle. Interestingly, Gemini-2.5-Pro is the only model that defies this trend, achieving a higher score in Reasoning (58.1) than in Perception (54.2). This exceptional result demonstrates that state-of-the-art models are beginning to overcome the reasoning bottleneck, showcasing advanced cognitive abilities that are on par with, or even exceed, their perceptual skills. The design of our dataset successfully magnifies this critical capability gap between the SOTA and other models.
@@
Baichuan-Omni-1.5 15.4 8.2 5.33 25.2 9 7.25 5.75 18.9 10.28 7.44 5.7 20.16
MiniCPM-O-2.6 20.0 6.2 11.33 29.6 9.05 9.55 8.02 22.3 11.24 8.88 8.43 23.76
Qwen-2.5-Omni-3B 19.8 12.2 5.33 33.6 10.7 7.2 8.86 22.55 12.52 8.2 8.42 24.76
Ming-lite-Omni-1.5 19.6 12.4 4.67 33.4 10.9 8.4 7.92 23.5 12.64 9.2 7.52 25.48
Qwen-2.5-Omni-7B 20.2 15.0 12.0 38.8 12.15 8.99 7.83 24.95 13.76 10.2 8.35 27.72
Qwen-3-Omni-30B-A3B 25.0 22.8 20.0 53.8 16.35 12.01 10.19 32.9 18.08 14.18 11.42 37.08
Gemini-2.0-Flash 25.2 19.4 14.67 49.0 15.5 14.05 13.02 35.95 17.44 15.12 13.22 38.56
Gemini-2.5-Flash 31.6 22.6 12.0 57.8 18.35 17.35 16.42 44.4 21.0 18.4 15.87 47.08
Gemini-2.5-Pro 25.6 22.2 21.33 54.2 22.9 25.05 19.43 58.1 23.44 24.48 19.67 57.32
@@ 4.4.2 Dataset Compression
The baseline data set consists of  8000 samples including 18 open-source benchmarks(e.g. MathVista and MMAU, details see Appendix.A) and 20 models evaluation results on them, which split into 12/8 on models as training/test set. Kmeans++(Arthur and Vassilvitskii, 2007) is used to cluster with K=48. To eliminate the random factor, we conduct 5-fold settings and evaluate 10 times on each setting.
To verify the compression performance, we define these metrics: Spearman’s Rank Correlation Coefficient(SRCC) for ranking consistency, Pearson’s Linear Correlation Coefficient(PLCC) for linear value consistency, Root Mean Square Error(RMSE) for numerical precision, Margin of Error(MoE) for quantifying estimation uncertainty, and Confidence Interval Coverage(CIC) for statistical reliability.
The experimental result is shown in Figure.9. At a 10% sampling rate, our method achieved excellent results on test-set. Both SRCC and PLCC exceeded 0.98, indicating near-perfect preservation of ranking and value relationships. The RMSE was below 0.02 with a corresponding MoE of 0.024; together, these values signify high numerical precision and a tight estimation range. Furthermore, the CIC was approximately 95%, confirming the statistical unbiasedness of the sample.
@@ 4.4.3 Benchmark Comparison
To ensure the quality of dataset, we conduct quality check on 10%-20% random samples in each benchmarks. As shown in Table.1, UNO-Bench has 100% accuracy on omni-modal dataset while 98% questions requires cross-modality to solve. It shows the highest quality among existing omni model benchmarks.
An effective benchmark must provide both a clear performance ladder and a meaningful difficulty range, and our UNO-Bench is engineered to deliver on both fronts as shown in Figure.10. It excels in discriminability, establishing substantial and remarkably linear intervals of ~10-12 points between adjacent models. This superior discriminability comes from a well-calibrated difficulty. UNO-Bench creates a vast 31.9 point performance gap between the top and bottom models, effectively separating their capabilities. This approach avoids the pitfall of being universally difficult, a problem seen in AV-Odyssey where all models are compressed into a narrow, low-scoring band (34-45). By combining a structured performance ladder with a balanced challenge, UNO-Bench serves as a more reliable and insightful tool for gauging genuine progress in the field.
@@ 5 Conclusion
1. With unified data framework in UNO-Bench, we found that the multimodal capability may not simply be a linear superposition of uni-modalities, but rather follows a significant multiplicative relationship. The evaluation results show that it manifests as a bottleneck effect on weak models, while exhibiting synergistic promotion on strong models.
@@
2. Both uni-modal and omni-modal understanding capability of the Gemini series far surpasses existing open-source omni models. The Gemini-2.5-Pro shows comparable perception capability with human experts but still has a performance gap in reasoning aspect.
@@
3. Besides better dataset quality and evaluation efficiency, UNO-Bench can provide sufficient metric discriminability and a progressive difficulty scale to drive model capability growth.
@@
In the future work, we will extend the dataset’s scale by human-in-the-loop automated pipeline and hold a private test set to avoid hacking. The ability coverage also needs to extend to more difficult reasoning tasks like STEM and code. At the same time, the relationship among cross-modals understandings and how to improve them are still exciting problem to explore. Furthermore, our compositional law has only been validated on UNO-Bench with uniformly distributed capabilities. Whether this law still holds under different capability distributions remains to be explored.
Acknowledgement
We hereby express our appreciation to the LongCat Team EVA Committee for their valuable assistance, guidance, and suggestions throughout the course of this work.
Appendix A Benchmarks Utilized in Dataset Compression
To construct our compressed datasets, we utilized a variety of benchmarks for both visual and audio modalities. For the visual component, we curated data from 15 public and several in-house benchmarks that assess a range of capabilities, including general visual question answering, document and chart comprehension, STEM/scientific reasoning, and multi-image understanding. For the audio component, we used 3 audio question answering benchmarks. The detailed composition of the resulting uni-modal dataset is presented in Figure.11.
Figure 11: The distribution of the uni-modal benchmarks in UNO-Bench. In addition to publicly available benchmarks, we incorporated several in-house benchmarks both before and after compression to ensure the reasonableness of the data distribution.
  • General visual question answering, RealWorldQAxAI (2023), MMEChaoyou et al. (2023), SeedBenchLi et al. (2023).
  • Document and chart understanding, OCRBench Liu et al. (2024d), FoxLiu et al. (2024a), DocLocal4kYe et al. (2023).
  • Stem & reasoning, MMMUYue et al. (2024), MMMU-ProYue et al. (2025), CMMMUGe et al. (2024), MathVistaLu et al. (2024b), MathVisionWang et al. (2024b), ScienceVistaTeam et al. (2025), GMAI-MMBenchYe et al. (2024).
  • Multi-image Understanding, ReMiKazemi et al. (2024), MuirBenchWang et al. (2024a).
  • Audio question answering, MMAUSakshi et al. (2025), MMSUWang et al. (2025), SDQAFaisal et al. (2021).
Appendix B Rigorous Derivation of the Compositional Law
Defining the performance gain as $\mathcal{P}^{\prime}_{\text{Omni}}=\mathcal{P}_{\text{Omni}}-b$. From Eq. 1, we have:
$\mathcal{P}^{\prime}_{\text{Omni}}(\mathcal{P}_{\text{A}},\mathcal{P}_{\text{V}})=f_{\text{A}}(\mathcal{P}_{\text{A}})+f_{\text{V}}(\mathcal{P}_{\text{V}})+f_{\text{I}}(\mathcal{P}_{\text{A}},\mathcal{P}_{\text{V}})$ (5)
Due to the high quality of our benchmark, where a task is unsolvable if either modality is absent, causing the performance to drop to its baseline (e.g. random guessing). we can have a strict boundary condition:
$\mathcal{P^{\prime}}_{\text{Omni}}(\mathcal{P}_{\text{A}},0)=0$ and (6)
$\mathcal{P^{\prime}}_{\text{Omni}}(0,\mathcal{P}_{\text{V}})=0$ and
$\mathcal{P^{\prime}}_{\text{Omni}}(0,0)=0$
Applying the boundary condition of Eq. 6 to Eq. 5, we find that the gain is a second-order mixed difference of $f_{\text{I}}$:
$\begin{split}\mathcal{P}^{\prime}_{\text{Omni}}(\mathcal{P}_{\text{A}},\mathcal{P}_{\text{V}})=&f_{\text{I}}(\mathcal{P}_{\text{A}},\mathcal{P}_{\text{V}})-f_{\text{I}}(\mathcal{P}_{\text{A}},0)\\ &-f_{\text{I}}(0,\mathcal{P}_{\text{V}})+f_{\text{I}}(0,0)\end{split}$ (7)
Substituting the Taylor series of $f_{\text{I}}$ around $(0,0)$ into Eq. 7, the performance gain is thus exactly equal to the sum of all pure interaction terms from $f_{\text{I}}$:
$\begin{split}\mathcal{P}^{\prime}_{\text{Omni}}(\mathcal{P}_{\text{A}},\mathcal{P}_{\text{V}})&=\sum_{i\geq 1,j\geq 1}c_{ij}\mathcal{P}_{\text{A}}^{i}\mathcal{P}_{\text{V}}^{j}\end{split}$ (8)
where the coefficients $c_{ij}$ are constants derived from the partial derivatives of $f_{\text{I}}$ at the origin. For sufficiently small uni-modal performances, we can approximate this series by its leading-order term:
$\mathcal{P}^{\prime}_{\text{Omni}}(\mathcal{P}_{\text{A}},\mathcal{P}_{\text{V}})\approx c_{11}\mathcal{P}_{\text{A}}\mathcal{P}_{\text{V}}$ (9)
This result strongly motivates modeling the interaction with the general multiplicative Cobb-Douglas form. Re-introducing the baseline $b$ yields our final Compositional Law:
$\mathcal{P}_{\text{Omni}}=C\cdot\mathcal{P}_{\text{A}}^{\alpha}\mathcal{P}_{\text{V}}^{\beta}+b$ (10)
Appendix C Model Selection for the Compositional Law
To validate our choice of the Compositional Law, we compared its performance against several alternative models. The fitting results for all candidate models on our 9-model dataset are summarized in Table.7.
Table 7: Fitting results for all candidate models. While more complex models achieve higher fitting scores ($R^{2}$), their parameters lack physical interpretability (e.g., negative exponents), indicating severe overfitting on our small dataset. Our chosen Symmetric Power Law offers the best balance of a high $R^{2}$ value and theoretical soundness.
Model Name $R^{2}$ RMSE Fitted Equation
Generalized Power Law 0.999 0.005 $P_{Omni}\approx 1.33\cdot P_{Audio}^{-1.59}\cdot P_{Visual}^{5.09}+0.24$
Linear Interaction 0.995 0.010 $P_{Omni}\approx 0.97-2.01P_{Audio}-0.59P_{Visual}+2.85(P_{Audio}\times P_{Visual})$
Weighted Sum Power Law 0.995 0.010 $P_{Omni}\approx 1.19\cdot(-0.20P_{Audio}+1.20P_{Visual})^{3.83}+0.24$
Symmetric Power Law 0.976 0.022 $\boldsymbol{P_{Omni}\approx 1.03\cdot(P_{Audio}\times P_{Visual})^{2.19}+0.24}$
Simple Linear 0.945 0.033 $P_{Omni}\approx-0.15-0.37P_{Audio}+1.43P_{Visual}$
As shown in Table.7, more complex models like the ‘Generalized Power Law’ achieve a near-perfect fit on the training data. However, this superior performance is misleading. These models yield parameters that are physically implausible, such as negative exponents (e.g., $P_{Audio}^{-1.59}$) or negative weights. Such parameters would illogically imply that improving a model’s uni-modal capability could degrade its omni-modal performance. This is a classic symptom of overfitting, where a model with high capacity learns the noise in a small dataset rather than a generalizable underlying trend.
In contrast, our proposed Symmetric Power Law provides an excellent fit ($R^{2}=0.976$) while maintaining theoretical coherence. All its parameters are positive and have clear interpretations: a super-linear synergy ($\alpha=2.19>1$) between modalities, a positive scaling factor ($C=1.03$), and a reasonable baseline score ($b=0.24$). Following the principle of Occam’s Razor, we select this model as it offers the most parsimonious, robust, and interpretable explanation for the observed phenomenon.
Interestingly, while the parameters from the overfitted models are invalid, they consistently suggest a stronger influence from the visual modality (e.g., the large positive exponent for $P_{Visual}$ in the ‘Generalized Power Law‘). This hints that while our symmetric law captures the primary collaborative effect, the visual component may play a slightly more dominant role, a direction for future investigation.
Appendix D General Scoring Model
One of the critical way to improve accuracy is to group questions into finer types and define appropriate criteria for each types. We define six different question types in Figure.12.
Figure 12: Definition of finer question types for general scoring model.
Appendix E Model Ability Taxonomy
This section will provide specific definitions for each ability item and present examples of various task types.
The specific model abilities and task types of the Perception dimension can be seen in Figure.13, and the Reasoning dimension can be seen in Figure.14.
Figure 13: Definition of the Perception Dimension.
Figure 14: Definition of the Reasoning Dimension.
Specific examples are provided for each model ability. Examples of Perception ability including Object Perception, Spatial Perception, Cross-Modal Alignment, Attribute Perception, Scenario Perception, Cross-Modal Conversion and Semantic Understanding can be seen in Figure.15. Examples of Reasoning ability including Complex Reasoning, Temporal Reasoning, Spatial Reasoning, Life Reasoning, STEM Reasoning and Code can be respectively seen in Figure.16.
Figure 15: Example of each ability in perception dimension.
Figure 16: Example of each ability in reason dimension.
Appendix F Omni-modal Solubility and Uni-modal Solubility Cases
Omni-modal solubility means that both visual and audio information are required to answer the question. While uni-modal solubility means that only visual or audio information is required to answer the question.
There are two uni-modal solubility cases. One is when audio and video present the same amount of information, and only one modality is needed to solve the problem, as shown in Figure.17(b). The other is when one modality provides useful information, while the other modality is unhelpful, as shown in Figure.17(c). The omni-modal solubility case can be seen in Figure.17(a).
Figure 17: Comparison of the problem solvable by each modality.